Y=C+I+G ,C=120+0.8Y , I=100+0.1Y ,G=300



Using cramer's rule find Y,C and I

1
Expert's answer
2022-02-03T11:52:09-0500

Y=C+I+G, C=120+0.8Y, I=100+0.1Y, G=300

Using Cramer’s rule find Y, C and I

This rule provides for methods of solving linear equations through use of determinants.

Y=120+0.8Y+100+0.1Y+300

Y-0.8Y-0.1Y=120+100+300

0.1Y=520

0.1Y/0.1=520/0.1

Y=5200

C=120+0.8Y

C=120+0.8*5200

C=120+4160

C=4280

I=100+0.1Y

I=100+0.1*5200

I=100+520

I=620